# Scaling Equitable Reflection Assessment in Education via Large Language Models and Role-Based Feedback Agents

### v1

@@ Introduction
We address these gaps with a theory-grounded, multi-agent pipeline [^1] Code, prompts, and anonymized data are available at https://github.com/CharlieChenyuZhang/equitable-reflection-assessment. that (1) produces robust, equitable rubric scores and (2) delivers dialogic feedback aligned with learning-science theory. By integrating advances in natural language processing with principles from the learning sciences, our approach provides formative feedback that is accurate and scalable while remaining pedagogically grounded and attentive to equity.
@@ 3. RQ3:
    Evaluator Agent applies the four-dimension rubric (Table 1) to the raw reflection text and produces structured scores along with brief reasoning. This agent grounds the workflow in an explicit, auditable rubric that instructors already use.
    Listing 1: Structured JSON returned by the Evaluator Agent
    1{
    2 "concept_understanding_score": 0,
    3 "real_world_application_score": 0,
    4 "reflection_questions_score": 0,
    5 "clarity_communication_score": 0,
    6 "reasoning": {},
    7 "areas_for_improvement": []
    8}
@@
Open Data.
An anonymized Google Sheet containing raw human scores, Likert ratings, and reflection IDs, alongside evaluation scripts and prompts, accompanies the public code release at https://github.com/CharlieChenyuZhang/equitable-reflection-assessment.
Table 1: Four-dimension grading rubric (0–3 points per dimension).
@@
A single pass of the grading and feedback pipeline using gpt-4o-mini-2024-07-18 consumes 1 2161\,216 input tokens and 2 2832\,283 output tokens, for a total of 3 4993\,499 tokens. At the model’s 18 July 2024 pricing (\$0.15 per million input tokens and \$0.60 per million output tokens), the cost per reflection is \$ $\mathdollar\,\mathbf{1.55\times 10^{-3}}$$221\,216\times 0.15/10^{6}+2\,283\times 0.60/10^{6}=0.001552$2. Grading and providing feedback for the 28-student cohort across three sessions ($28\times 3=84$ reflections) therefore costs \$ $\mathdollar\,0.13$. Applying the same process to all twelve weekly reflections ($28\times 12$) yields a total cost of \$ $\mathdollar\,0.52$.
@@ Conclusion
Reproducibility Checklist
  1. 1.
    Includes a conceptual outline and/or pseudocode description of AI methods introduced — Yes.
  2. 2.
    Clearly delineates statements that are opinions, hypotheses, and speculation from objective facts and results — Yes.
  3. 3.
    Provides well-marked pedagogical references for less-familiar readers — Yes.
  4. 4.
    Does this paper make theoretical contributions? — No
  5. 5.
    Does this paper rely on one or more datasets? — Yes
    • Motivation for choosing the dataset(s) is provided — Yes
    • All novel datasets are included in a data appendix — Yes
    • Novel datasets will be publicly released under a research-friendly licence — Yes
    • Datasets from prior literature are properly cited — NA
    • All cited external datasets are publicly available — NA
    • Non-public datasets are described and justified — NA
  6. 6.
    Does this paper include computational experiments? — Yes
    • Hyper-parameter search ranges and selection criteria are reported — Partial. (Fixed temperature; other defaults noted in §Methodology.)
    • Pre-processing code is included in the appendix — Yes
    • All code for running and analyzing experiments is in a code appendix — Yes
    • Code will be released publicly under a research-friendly licence — Yes
    • New-method source code is commented and references paper steps — Yes
    • Random-seed handling is described to enable replication — Partial. (Deterministic seed for evaluation scripts; LLM API noted.)
    • Computing infrastructure (hardware & software) is specified — Yes.
    • Evaluation metrics are formally defined and motivated — Yes
    • Number of algorithm runs per reported result is stated — Yes.
    • Experimental analysis reports measures of variation or confidence — Yes.
    • Statistical significance of performance differences is tested — Yes
    • Final hyper-parameters for each model/algorithm are listed — Partial. (LLM temperature, max-tokens, and role prompts given.)

### v2

@@ Abstract
Code, Prompts, and Anonymized Datahttps://github.com/CharlieChenyuZhang/equitable-reflection-assessment
@@ Introduction
We address these gaps with a theory-grounded, multi-agent pipeline that (1) produces robust, equitable rubric scores and (2) delivers dialogic feedback aligned with learning-science theory. By integrating advances in natural language processing with principles from the learning sciences, our approach provides formative feedback that is accurate and scalable while remaining pedagogically grounded and attentive to equity.
@@ 3. RQ3:
Student Reflection Evaluator Agent Equity Monitor Agent Metacognitive Agent Aggregator Agent Reflexion Agent
@@
    Evaluator Agent applies the four dimension rubric to the raw reflection text and returns a structured JSON object. The output includes (i) an integer score from 0 to 3 for each rubric dimension, (ii) a brief natural language explanation of the score in a reasoning field, and (iii) an areas_for_improvement list that highlights specific suggestions for the learner. This structured format provides an explicit, auditable record of how each score is assigned and why.
@@
Table 1: Four-dimension grading rubric (0–3 points per dimension).
@@
[…]